OpenStudy (natsent06):

What value of n will make (a + b) × n = 1 a TRUE statement?

OpenStudy (natsent06):

@iGreen

OpenStudy (ahsome):

Impossible without info on a and b

OpenStudy (natsent06):

@Ahsome A) (a + b) B) (a - b) C) (2a + 2b) D) 1/(a + b)

OpenStudy (ahsome):

D. Since: \[\frac{1}{(a+b)}*(a+b)=1\]

OpenStudy (natsent06):

@iGreen help

OpenStudy (natsent06):

Find the value of n which will make (3/4) × n = 12/16 a TRUE statement. A) 16/16 B) 9/9 C) 24/24 D) 4/4

OpenStudy (natsent06):

its like the other question

OpenStudy (ahsome):

All of them, because they all equal to 1

OpenStudy (natsent06):

well i can only pick one

OpenStudy (ahsome):

But 16/16=1, 9/9=1, 24/24=1 and 4/4=1. Not sure now :(

OpenStudy (natsent06):

@iGreen !!!!!

OpenStudy (igreen):

Idk..sorry.. :l

OpenStudy (natsent06):

ARGHSRTNBSVIUHM GIUHT

OpenStudy (natsent06):

*rages*